ganeshie8 (ganeshie8):

How to split |sinx| into two parts such that one part is fixed and the other varying?

ILovePuppiesLol (ilovepuppieslol):

whats a sinx

OpenStudy (blueyoshi):

Why are 1000 people helping this person and not one will help me...

OpenStudy (anonymous):

could someone help me? No one is one my question, and this one only needs 2.

OpenStudy (anonymous):

*on

ganeshie8 (ganeshie8):

The usual trig sine function. I need to split it something like below |sin x| = C + f(x)

ILovePuppiesLol (ilovepuppieslol):

aw i don't know this then :(

Parth (parthkohli):

f(x) couldn't be a neat function because it's not gonna be differentiable.

OpenStudy (johnweldon1993):

^My thinking as well >.<

OpenStudy (blueyoshi):

Seriously... I need a lot of help and there are like 10 people hleping this guy and not one will help me with a problem much simpler than this..

ganeshie8 (ganeshie8):

That's okay I just need to prove that it is possible to split any "pulsating dc" into sime fixed DC plus something varying

Parth (parthkohli):

Is there a good-looking answer for this? I highly doubt it though since \(|\sin x| - c\) isn't gonna be good-looking :|

Parth (parthkohli):

oh, then that fixed value could be the mean-position and the other could be an \(A \sin (\omega t + \phi)\)-type function which oscillates about that mean value

ganeshie8 (ganeshie8):

Awesome So just take the fixed DC to be the average value of the pulsating DC?

OpenStudy (anonymous):

these people are jerks

ganeshie8 (ganeshie8):

I'm on mobile Not seeing drawing tool n medal buttons

OpenStudy (anonymous):

not a single one is available to help us, but they'll all flock over to help someone with the "Professor" title.

Parth (parthkohli):

Yeah, whatever an "average" value is, I'm sure that it turns out to be the value it's oscillating about. I think you're right.

ganeshie8 (ganeshie8):

OK that's the rough idea. Still need to prove though

Parth (parthkohli):

I think that's what Fourier's Theorem is though - that any periodic thing can be expressed as the sum of sines added to a constant.

Parth (parthkohli):

@Kainui should know more about the proof.

ganeshie8 (ganeshie8):

Ahh I see. There is a simple electric component that does this splitting in no time

ganeshie8 (ganeshie8):

It is called choke input filter I'm looking for some mathematical treatment of it

ganeshie8 (ganeshie8):

The fixed dc value in that link is indeed the average value of |sinx| which is 2/pi But that link doesn't explain much about the varying part

OpenStudy (kainui):

I was gonna say, looks like full-wave rectifier. Here's a great video from a channel I found about a year ago describing this, pretty cool stuff imo https://www.youtube.com/watch?v=cyhzpFqXwdA Here in your picture, they're taking AC current and turning it into DC current with some sorta filter, and that filter apparently seems to make it oscillate slightly around the average height. Just for fun, here's the calculation of the average (which is the same value given on the right there, where they've written \(V_{dc} = \frac{2}{\pi} V_{L max}\). V is the amplitude, a constant. \[\frac{1}{2 \pi}\int_0^{2 \pi} V|\sin x| dx = \frac{V}{\pi} \int_0^\pi \sin x dx = \frac{2V}{\pi}\]

ganeshie8 (ganeshie8):

Exactly! so the choke input filter gives the output at average value In that youtube video the filter gives the output at Vmax value !